OpenStudy (anonymous):

At the grocery store, 5 pints of ice cream cost $13.10. How much would 35 pints of ice cream cost?

OpenStudy (ghazi):

5 costs $13.10 1 would cost = 13.10/5 =$2.62 therefore 35 would cost 35*2.62=$ 91.7

OpenStudy (anonymous):

@ghazi thank you so much.

OpenStudy (ghazi):

:D :)

OpenStudy (anonymous):

@ghazi can you help me with one more?

OpenStudy (ghazi):

sure :)

OpenStudy (anonymous):

Simplify: (-12)*2

OpenStudy (ghazi):

its just product of -12*-12= 144 because -*-=+

OpenStudy (anonymous):

thanks a lot!

OpenStudy (ghazi):

:) YW